S. Mobile phone service providers (e.g., Verizon, AT&T, T-Mobile) surveyed 100 households in each of the cities of New Orleans, Baton Rouge, and Lafayette. The results households, 80 households, and 65 households, respectively, in the three cities sti subscribed to land-line telephone services. Use the Chi-square analysis to test e that the household ownership of land-lines in each of the three cities Conduct the test at the 1% level of significance (0.01). showed that 85 is equal to 70 percent. 6. Young-Woodley Baking Company was interested in evaluating the packages for its cookie products. The firm selected 6 of its stores with sim of cookies and randomly split them into two groups of three stores each. one group was stocked only with the round designer packages, while the othe roupsls to stocked only with square packages. All the stores were monitored for two certain that the packages were properly sales effect of two different ilar monthly sales of stores of stores was weeks to make displayed and that no stock-outs occurred. After the two-week period of the experiment, the following results were obtained. variance (ANOVA) test to see if the two different types of boxes made any dfere sales of the cookies. Conduct the test at the 5% level of significance (0.05) Store A Store V Store C 10 12 14 Store E Store W Store D Square Package 14 16 18 Total Sales Both Packages 24 28 32 Total Mean 36 12 48 16 84 14
Answer #1

null hypothesis: Ho: each city has ownership of landlines equal to 70%.

alternate hypothesis:Ha: at least one city has ownership different from 70%

Observed Oi New Orleans Baton Rouge Lafayette Total
Land lines 85 80 65 230
no land line 15 20 35 70
Total 100 100 100 300
Expected Ei=Σrow*Σcolumn/Σtotal New Orleans Baton Rouge Lafayette Total
Land lines 70.0 70.0 70.0 210
no land line 30.0 30.0 30.0 90
Total 100 100 100 300
chi square    χ2 =(Oi-Ei)2/Ei New Orleans Baton Rouge Lafayette Total
Land lines 3.2143 1.4286 0.3571 5.000
no land line 7.5000 3.3333 0.8333 11.667
Total 10.714 4.762 1.190 16.6666

test statsitic =16.67

degree of freedom(df) =(rows-1)*(columns-1)=

2

for 2 df and 0.01 level of signifcance critical region       χ2= 9.210

Decision:as test statitic is in rejection region we reject null hypothesis

Conclusion: we have sufficient evidence to conclude that at least one city has ownership different from 70%

------------------------

only one ques can be answered per post as HOMEWORKLIB RULES guidelines.